# Multi-Parameter Rating Methodology for Systematic Comparison of Theoretical Models with Experimental Data in Heavy-Ion Physics

### v1

@@ 4.1 Rating Calculation Example: PYTHIA8 for Mesons
$R$ $=0.45\,Q_{w}+0.30\,Q_{g}+0.20\,Q_{\min}$
$\quad-0.05\,\tfrac{\sigma}{300}$
$=0.45(963)+0.30(937)+0.20(823)$
$\quad-0.05\cdot\tfrac{52.9}{300}\approx\mathbf{879}$
@@ 5 Advantages Over Traditional $\chi^{2}$/ndf
The multi-parameter rating methodology offers six key advantages:
@@
(3) Robustness: Geometric mean and minimum aggregators prevent compensation cheating where poor agreement in one region is artificially balanced by overfitting another. Model with $Q_{\text{weighted}}=850$ but $Q_{\text{minimum}}=200$ receives lower rating $R$ than model with uniform $Q_{i}=750$ everywhere.
(4) Transparency: Spider plots (Fig. 1) and heatmaps (Fig. 2) enable instant identification of model strengths/weaknesses without detailed statistical analysis. Non-experts can immediately see that PYTHIA8 excels at jets while EPOS dominates thermal regime.
@@ 6 Development Perspectives
Future applications include:
@@
(7) Precise model-by-model $Q_{i}$ calculation: Full integration with Monte Carlo event generators (EPOS-LHC, PYTHIA8, AMPT) to compute exact $Q_{i}$ values across all observables and zones via direct $\chi^{2}_{i}/\nu_{i}$ statistics. Current work demonstrates methodology with representative estimates; complete implementation requires systematic model output comparison with experimental data point-by-point.
@@ 7 Conclusions
(1) Seven-zone division captures all essential mechanisms from thermal freeze-out ($p_{T}<0.8$ GeV/c) through hydrodynamic flow, Cronin enhancement, critical transition, hard fragmentation, to asymptotic pQCD ($p_{T}>10$ GeV/c). Zone 4 (2.5–4.0 GeV/c) emerges as most discriminating due to soft-hard interplay.
(2) Integrated rating $R$ combines Bayesian logic (geometric mean penalizes non-uniformity) with multi-criteria decision making (weighted average emphasizes important zones, minimum provides conservative bound). Prevents compensation artifacts inherent in global $\chi^{2}$/ndf.
(3) Visualization tools (spider plots in Fig. 1, heatmaps in Fig. 2, scatter diagrams in Fig. 3) enable immediate identification of model complementarity: hydrodynamic models excel at low $p_{T}$, pQCD generators at high $p_{T}$, only microscopic transport with coalescence succeeds for baryon anomaly as seen in Fig. 8.
(4) Application to real ALICE data for $K^{0}_{S}$ at 5.02 TeV demonstrates PYTHIA8 best for mesons ($R=879$) via nuclear PDFs, while the baryon anomaly pattern in Fig. 8 highlights the necessity of coalescence mechanisms for baryons. Analysis for $\Lambda$ baryons continues and requires full simulation of models with coalescence mechanisms (such as PHSD).
(5) Methodology ready for practical use: transparent (all calculations reproducible from Tables 14), scalable (extends to $v_{n}$, HBT, heavy flavor), automated (database schema enables batch processing). Future integration into standard analysis workflows will accelerate progress in theoretical QGP description.
@@
Acknowledgments
We acknowledge the organizers of the 2nd Workshop on Nuclear Physics and High-Energy Physics: Theory, Experiment, and Applications (Kyiv, Ukraine, 2026) for the opportunity to present preliminary results of this work [22].

### v2

@@ 4.1 Rating Calculation Example: PYTHIA8 for Mesons
$R$ $=0.45\,Q_{w}+0.30\,Q_{g}+0.20\,Q_{\min}-0.05\cdot\tfrac{\sigma}{\sigma_{\max}(\mathbf{w})}$
$=0.45(963)+0.30(937)+0.20(823)-0.05\cdot\tfrac{52.9}{450}$
$=433.4+281.1+164.6-0.009\approx\mathbf{879}$
@@ 5 Advantages Over Traditional $\chi^{2}$/ndf
The multi-parameter rating methodology offers seven key advantages.
@@
(3) Robustness: Geometric mean and minimum aggregators prevent compensation cheating where poor agreement in one region is artificially balanced by overfitting another. A model with $Q_{\text{weighted}}=850$ but $Q_{\text{minimum}}=200$ receives lower rating $R$ than a model with uniform $Q_{i}=750$ everywhere.
(4) Transparency: Spider plots (figure 1) and heatmaps (figure 2) enable instant identification of model strengths/weaknesses without detailed statistical analysis. Non-experts can immediately see that PYTHIA8 excels at jets while EPOS dominates the thermal regime.
@@
(7) Cross-system comparability: Physically motivated zone boundaries encode universal QCD properties — chemical freeze-out temperature, onset of radial flow, Cronin threshold, soft-to-hard transition — that are independent of collision system, beam energy, and detector acceptance. The same partition $\{0.8,\,1.6,\,2.5,\,4.0,\,6.0,\,10.0\}$ GeV/$c$ applies directly to p-Pb, Pb-Pb, Xe-Xe, and future O-O data, enabling genuine inter-system model comparisons. Data-driven boundaries, optimized per dataset, are incommensurable across experiments: a model’s rating on ALICE p-Pb data cannot be directly compared with its rating on CMS Pb-Pb data, undermining the goal of a community-wide evaluation framework.
@@ 6 Development Perspectives
Future applications include the following directions.
@@
(7) Precise model-by-model $Q_{i}$ calculation: Full integration with Monte Carlo event generators (EPOS-LHC, PYTHIA8, AMPT) to compute exact $Q_{i}$ values across all observables and zones via direct $\chi^{2}_{i}/\nu_{i}$ statistics. The present work demonstrates the complete methodology on digitized model predictions for $K^{0}_{S}$ (table 9) and on a synthetic benchmark for $\Lambda/K^{0}_{S}$; systematic point-by-point comparison for all eight models across all seven zones is reserved for a dedicated follow-up study.
@@ 7 Conclusions
(1) Seven-zone division captures all essential mechanisms from thermal freeze-out ($p_{T}<0.8$ GeV/$c$) through hydrodynamic flow, Cronin enhancement, critical transition, hard fragmentation, to asymptotic pQCD ($p_{T}>10$ GeV/$c$). Zone 4 (2.5–4.0 GeV/$c$) emerges as most discriminating due to soft-hard interplay. The physically motivated zone-division strategy is demonstrated to outperform data-driven and hybrid alternatives in reproducibility and physical interpretability (section 2.1): boundary significance tests on CMS data independently confirm 4 of 6 physical boundaries at $p<0.05$, while none of the three data-driven K=9K{=}9 candidate boundaries pass the same test independently. Crucially, data-driven AIC/BIC selection is shown to be mathematically equivalent to a dataset-specific reweighted $\chi^{2}$/ndf, recovering the very averaging problem the methodology is designed to solve; physically motivated boundaries, encoding universal QCD properties, are the only choice that guarantees cross-system comparability of model ratings.
(2) Integrated rating $R$ combines Bayesian logic (geometric mean penalizes non-uniformity) with multi-criteria decision making (weighted average emphasizes important zones, minimum provides conservative bound), and includes a bounded dispersion correction. Prevents compensation artifacts inherent in global $\chi^{2}$/ndf. The three primary aggregator coefficients $(0.45,0.30,0.20)$ form a geometric sequence with ratio $2/3$, reflecting diminishing statistical information from mean to geometric mean to infimum; the bounded dispersion correction $\bigl(-0.05\,\sigma/\sigma_{\max}(\mathbf{w})\bigr)$ uses a weight-normalised denominator $\sigma_{\max}(\mathbf{w})\approx 450$ so that the four aggregation coefficients form a strict partition of unity and the correction is bounded to [0.05, 0][-0.05,\,0] rating points, while $Q_{\text{geometric}}$ carries the substantive non-uniformity penalty. The robustness of the aggregation is quantitatively confirmed: $\pm 20\%$ variations in all four coefficients produce $\Delta R<2\%$ with zero rank changes across all evaluated models (table 5).
(3) Visualization tools (spider plots in figure 1, heatmaps in figure 2, scatter diagrams in figure 3) enable immediate identification of model complementarity: hydrodynamic models excel at low $p_{T}$, pQCD generators at high $p_{T}$, only microscopic transport with coalescence succeeds for baryon anomaly as seen in figure 8.
(4) Application to real ALICE data for $K^{0}_{S}$ at 5.02 TeV demonstrates PYTHIA8 best for mesons ($R=879$) via nuclear PDFs. The baryon anomaly peak in zones 3–4 yields quality scores $Q_{i}\sim 950$$1000$ for models incorporating coalescence (PHSD base $R=893$), versus $Q_{i}\sim 400$$600$ for pure fragmentation generators — confirming the role of the $\Lambda/K^{0}_{S}$ ratio as the strongest model discriminator. The ranking for $\Lambda$ baryons is demonstrated here on a synthetic benchmark consistent with LHCb kinematics; a dedicated study with the full experimental dataset will provide final rankings for the baryon sector.
(5) Methodology ready for integration into standard analysis frameworks: transparent (all calculations reproducible from tables 16), scalable (extends to $v_{n}$, HBT, heavy flavor), and fully algorithmic (database schema enables batch processing of dozens of models). The framework is designed for direct incorporation into ALICE, CMS, and JETscape workflows without conceptual modification.
@@
Acknowledgements.
We acknowledge the organizers of the 2nd Workshop on Nuclear Physics and High-Energy Physics: Theory, Experiment, and Applications (Kyiv, Ukraine, 2026) for the opportunity to present preliminary results of this work indico2026 .